OpenStudy (anonymous):

PLEASE HELP Find the average rate of change of g(x) = -x^2 +4(x) +1 from x= 3 to x=6 . Simplify your answer as much as possible.

OpenStudy (anonymous):

g(6) = -6^2 + 4(6) +1 g(3) = -3^2 + 4(3) +1 then do I subtract the second equation from the first?

OpenStudy (anonymous):

@KendrickLamar2014

OpenStudy (kendricklamar2014):

I really dont know how to solve this :( maybe @Michele_Laino or @tkhunny can help you

OpenStudy (anonymous):

ok =) when I plugged in 6 g(6) = -11 when I plugged in 3 g(3) = 4 ( not sure If i'm on the right track though or if I literally need to subtract the equations)

OpenStudy (anonymous):

@satellite73

OpenStudy (kendricklamar2014):

@amistre64

OpenStudy (kendricklamar2014):

@paki @triciaal

OpenStudy (anonymous):

thank you =) I want to say it's -11-4 but I vaguely remember something about subtracting the equations or something slightly more complicated soo

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

@BloomLocke367

OpenStudy (anonymous):

I'm guessing I probably did wrong

OpenStudy (bloomlocke367):

I actually don't know.. sorry! I thought you meant with linear equations..

OpenStudy (anonymous):

it's ok thanks anyways =)

OpenStudy (triciaal):